OpenStudy (augustmae):

How would I solve the equation limit of 4x^2/x-2 while x tends to ∞? I can solve limits with numbers, but I wasn't sure what to do in this situation. Would I just eliminate?

OpenStudy (freckles):

is that 4x^2/(x-2) or 4x^2/x-2 ?

OpenStudy (freckles):

\[ \text{ I believe you meant } \lim_{x \rightarrow \infty} 4x^2/(x-2) \\ \text{ so let's look at } \lim_{x \rightarrow \infty} \frac{4x^2}{x-2}\]

OpenStudy (freckles):

Let me ask you if you know the following answers: \[\lim_{x \rightarrow \infty} 4x^2= ? \\ \lim_{x \rightarrow \infty}(x-2)=?\]

OpenStudy (augustmae):

lim if 4x^2 = ∞ lim of (x-2) = ∞

OpenStudy (freckles):

right do you think one function gets bigger faster

OpenStudy (freckles):

I'm talking about comparing f(x)=4x^2 to g(x)=x-2 and telling me if you think one takes off way faster in getting bigger

OpenStudy (augustmae):

I'm just guessing here, but would it be the function f(x) = 4x^2?

OpenStudy (augustmae):

because the function is squared?

OpenStudy (freckles):

yep you can also look at the graph there the 4x^2 is way bigger than the x-2 for really big values of x

OpenStudy (freckles):

that means you will be dividing really big numbers by other numbers that are less big

OpenStudy (freckles):

what do you think that means about the limit to your initial question?

OpenStudy (augustmae):

the limit will go to infinity

OpenStudy (freckles):

right if the 4x^2 was on bottom and the x-2 was on top than you would say 0 instead

OpenStudy (augustmae):

Ok. I get it now

OpenStudy (augustmae):

Thank You!

OpenStudy (freckles):

if you had the same degree though then you would need to do something differently

OpenStudy (freckles):

same degree means the top and bottom has the same largest exponent

OpenStudy (freckles):

for example \[\lim_{n \rightarrow \infty} \frac{4x^2}{x^2-2}=\frac{4}{1}=4 \\ \text{ I just stole the coefficients of } x^2 \text{ on top and on bottom }\] coefficient of 4x^2 is 4 coefficient of x^2 is 1

OpenStudy (freckles):

oops that was suppose to say x approaches infinity not n

OpenStudy (freckles):

same degree examples: \[\lim_{x \rightarrow \infty} \frac{4x^2}{x^2-1}=\frac{4}{1}=4 \\ \lim_{x \rightarrow \infty} \frac{4x^3}{x^2-x^3} =\frac{4}{-1}=-4 \\ \lim_{x \rightarrow \infty} \frac{5x-1}{6x-2}=\frac{5}{6} \]

OpenStudy (freckles):

bigger degree on top examples: \[\lim_{x \rightarrow \infty} \frac{4x^3}{x-1}=\infty \\ \lim_{x \rightarrow \infty} \frac{x^4-1}{4x^2+1}=\infty\]

OpenStudy (freckles):

bigger degree on bottom examples: \[\lim_{x \rightarrow \infty} \frac{4x}{2x^3-1}=0 \\ \lim_{x \rightarrow \infty} \frac{4x^3+5x^2+1}{8x^4}=0 \\ \lim_{x \rightarrow \infty} \frac{1}{x}=0\]

OpenStudy (freckles):

notice all of these examples I mentioned are when you have a polynomial over a polynomial and x is approaching infinity

OpenStudy (freckles):

oops one more thing about the bigger degree on top examples

OpenStudy (freckles):

I chose ones that gave you positive infinity

OpenStudy (freckles):

let me choose someones that would give you negative infinity instead

OpenStudy (augustmae):

I got the answer for my question, but how does an answer become 0 or a number besides infinity?

OpenStudy (augustmae):

I'm not sure I understand how you got the answers for the examples

OpenStudy (freckles):

add to this above to my section on the degree on top is bigger examples: \[\lim_{x \rightarrow \infty} \frac{4x^3}{x-1}=\infty \\ \lim_{x \rightarrow \infty} \frac{x^4-1}{4x^2+1}=\infty \\ \lim_{x \rightarrow \infty } \frac{-x^4-1}{4x^2+1}=-\infty \\ \lim_{x \rightarrow \infty} \frac{x^4-1}{-4x^2+1}= - \infty \\ \lim_{x \rightarrow \infty} \frac{-x^4-1}{-4x^2+1}=\infty\]

OpenStudy (freckles):

did you understand my very first examples before beginning the other same degree examples

OpenStudy (freckles):

this one: \[\lim_{n \rightarrow \infty} \frac{4x^2}{x^2-2}=\frac{4}{1}=4 \\ \text{ I just stole the coefficients of } x^2 \text{ on \top and on bottom } \] coefficient of 4x^2 is 4 coefficient of x^2 is 1

OpenStudy (augustmae):

yes

OpenStudy (augustmae):

I understood that one

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \frac{4x^2}{x^2-1}=\frac{4}{1}=4 \\ \lim_{x \rightarrow \infty} \frac{4x^3}{x^2-x^3} =\frac{4}{-1}=-4 \\ \lim_{x \rightarrow \infty} \frac{5x-1}{6x-2}=\frac{5}{6} \] which of these same degree examples did you understand if any

OpenStudy (augustmae):

the first one. Do you just take the coefficients?

OpenStudy (freckles):

the coefficients of the terms with the largest exponents on top and bottom this is only if the largest exponents match that you can do this

OpenStudy (freckles):

the first example deg(4x^2)=deg(x^2-1) so you can apply this and say the limit is 4/1 or 4 the second example deg(4x^3)=deg(-x^3+x^2) so you can apply this and say the limit is 4/(-1) or -4 the third example degree(5x-1)=deg(6x-2) so you can apply this and say the limit is 5/6

OpenStudy (freckles):

further information if you need it: deg(4x^2)=deg(x^2-1)=2 deg(4x^3)=deg(-x^3+x^2)=3 deg(5x-1)=deg(6x-2)=1

OpenStudy (freckles):

so these are the limits when you have a polynomial over a polynomial as x approaches infinity or negative infinity: if the deg(top)=deg(bottom) then the answer is going to come from the coefficients of the terms with largest exponents if the deg(top)>deg(bottom) then the answer is going to be either -infinity or +infinity if the deg(top)<deg(bottom) then the answer is going to be zero

OpenStudy (freckles):

deg( ) just means the degree of the polynomial inside the deg( )

OpenStudy (augustmae):

ok

OpenStudy (augustmae):

I think i get it

OpenStudy (freckles):

do you know this limit then? \[\lim_{x \rightarrow \infty} \frac{-9x^2+5x-1}{3x^2}\]

OpenStudy (augustmae):

3

OpenStudy (freckles):

almost

OpenStudy (freckles):

the coefficient of -9x^2 is -9

OpenStudy (freckles):

so your answer should be -9/3 and if reduced just -3

OpenStudy (augustmae):

ok

OpenStudy (augustmae):

how do I determine the degree of the polynomial again?

OpenStudy (freckles):

it is largest exponent for examples: deg(5x^3-4x+1)=3 deg(2x-1)=1 deg(5)=0 deg(6x^2-1)=2

OpenStudy (augustmae):

now I remember

OpenStudy (augustmae):

sorry

OpenStudy (freckles):

np all of this was probably learned years before you entered this calculus class I don't even think it takes very much time to forget

OpenStudy (augustmae):

I think I got it. Thank You!

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/CalcI/LimitsAtInfinityI.aspx this contains more examples and it goes into more detail using the following limit to evaluate other limits \[\lim_{x \rightarrow \infty} \frac{1}{x^r}=0 \text{ where } r \text{ is a positive rational number }\]

OpenStudy (freckles):

the above things i mentioned were shortcuts

OpenStudy (freckles):

the site there mentions the non short cut way is what i'm trying to say

OpenStudy (freckles):

i will be back later if you have more questions